A monkey is decending from the branch of a tree with constant acceleration. If the breaking strength is 75% of the weight of the monkey, the minimum acceleration with which monkey can slide down without breaking the branch is (a) g (b) 3g/4 (c) g/4 (d) g/2

Grade:11

1 Answers

Rituraj Tiwari
3 years ago
Let T be the tension in the branch of a tree when monkey is descending with acceleration a. Then mg - T = ma; T = 75% of weight of monkey ma = mg - (75/100) mg= (1/4) mg or a =g/4.